OpenStudy (anonymous):

of ten twenty-dollar bills, 2 are counterfeit. six bills are chosen at random. what is the probability that both counterfeit bills are chosen?

OpenStudy (anonymous):

\[\frac{1}{3}\] there are two ways of doing this

OpenStudy (anonymous):

the number of ways to choose 6 from 10 is \[\dbinom{10}{6}=210\] and there is only one way to choose both counterfeit bills and \[\dbinom{8}{4}=70\] ways to choose the other 4 so the answer is \[\frac{70}{210}=\frac{1}{3}\]

OpenStudy (anonymous):

second way is to say that the probability first is counterfeit is \[\frac{2}{10}\] and then the probability that the second is counterfeit given the first one is is \[\frac{1}{9}\] and then the other bills are all good. but there are \[\dbinom{6}{2}=15\] ways to arrange the counterfeit bills in the set of 6 chosen, so the probability is \[15\times \frac{2}{10}\times \frac{1}{9}=\frac{1}{3}\]